# Noise-resilient host for electron qubit operation up to 0.4 K

### v1

@@
A separate measurement on a different qubit Q2, presented in Supplementary V, also reveals strikingly thermal behavior in $T_{1}$ and thermal population but with a different $T_{1}$ in the low-temperature limit. It suggests a picture in which individual eNe qubits couple differently to their environment (perhaps through phonons [14, 9]) which sets the low temperature $T_{1}$. This motivates further studies of the microscopic limitations of $T_{1}$. At these qubit frequencies, $T_{1}$ drops to only about half the low-temperature value at 200 mK, facilitating higher temperature operation. In contrast, electron qubits in semiconductor materials are susceptible to phonon effects [25, 49], which can significantly degrade $T_{1}$ at elevated temperatures. Given these favorable scaling, we can also anticipate further improvement of the high-temperature performance of eNe qubits by engineering the electron trapping and detection structure to support higher qubit frequency operation [24].
@@
The temperature dependence of the coherence data, however, remains complex. $T_{2}^{*}$ and $T_{2}^{\rm{echo}}$ show that quasi-static and high-frequency noise components behave differently with temperature, as shown in Fig. Noise-resilient host for electron qubit operation up to 0.4 Kb. With increasing temperature, $T_{2}^{*}\propto T^{-0.74}$, indicating the quasi-static noise power scales almost linearly against temperature. Below 50 mK, $T_{2}^{\rm{echo}}$ approaches $2T_{1}$, suggesting that the single echo pulse still effectively mitigates the quasi-static noise that contributes to most of the dephasing. However, starting from 75 mK, $T_{2}^{\rm{echo}}$ scales with temperature as $\propto T^{-1.1}$, and quickly degrades from $2T_{1}$, indicating the rise of non-quasistatic dephasing noises at higher temperatures [23]. Despite that, Q1 maintains $>$ 1 µs $T_{2}^{\rm{echo}}$ at elevated temperatures up to 400 mK, showing eNe’s thermal-resilient coherence performance and potential for high-temperature operation.
@@
Fig. 5. : Temperature-dependent coherence of qubit Q1 at charge sweet-spot. a, $T_{1}$ (blue dots, data) versus mixing chamber temperature. The solid curve represents the predicted decay rate $T_{1}(T)=T_{1}(T=0)\cdot\text{tanh}(\hbar\omega_{\rm{q}}/2k_{B}T)$ where we use the measured value at 10 mK for $T_{1}(T=0)$. Extracted electron temperature (gray diamonds, data) versus MXC temperature based on the measured thermal population in the Supplementary. b, Decoherence time, $T_{2}^{*}$ (orange dots) and $T_{2}^{\rm{echo}}$ (green dots) versus the MXC temperature, with power-law fitting. Gray dots and curves show $2T_{1}$ for comparison. c, Extracted pure dephasing time $T_{\phi}$ as a function of the MXC temperature. Red dashed curve: Parameter-free calculation of the resonator-induced dephasing based on Q1 properties.
@@
We then extract the relation between the pure dephasing time ($T_{\phi}$) and MXC temperature, following the coherence decay formula described in the Methods. As shown in Fig. Noise-resilient host for electron qubit operation up to 0.4 Kc, $T_{\phi}$ begins to decrease significantly when the MXC temperature exceeds 100 mK. In this range, the measured data matches well with a parameter-free model accounting for the effect of resonator thermal photon on the qubit dephasing [24, 50]:
@@
where $\kappa$ is the resonator decay rate, and $\chi$ is the resonator’s dispersive shift (see Supplementary IV). $n_{\text{th}}$ represents the resonator thermal photon population, equal to $1/(e^{hf_{\rm{r}}/k_{\text{B}}T}-1)$. At higher temperatures, the degradation of qubit coherence is primarily due to thermal effects. Nevertheless, the deviation of the data from the model below 100 mK suggests the presence of other low-temperature dephasing mechanisms. Separate measurements on Q2 (Supplementary V) give similar results.
@@
The inconsistent spectral properties of eNe qubits indicate that the local neon profile plays a crucial role in the electron-trapping mechanisms. Variations in the neon surface profile, caused by liquid-phase solidification and substrate roughness, introduce randomness in the energy potential landscape that hosts the electron qubits. Additionally, the relative position of trapped electrons to nearby electrodes determines the frequency lever-arm and the qubit’s sensitivity to voltage fluctuations. To enhance consistency and control over individual qubit properties, refined neon growth method and gate-defined electron trapping mechanisms should be developed to mitigate the effects of neon surface randomness.
@@
Finally, the limiting factor for the non-radiative energy relaxation rate of eNe qubits needs further investigation. The non-monotonic variations in $T_{1}$ of qubits with different charge sweet-spot frequencies point to localized non-radiative decay channels. The presence of nearby charges may create a sparse bath of two-level fluctuators weakly coupled to the eNe qubit, whose density determines the transverse noise intensity [53, 51]. Electrical manipulations could be applied to bias or repeal those weakly coupled charges to reduce the relaxation rate [54]. With the development of improved electron loading and trapping methods, we foresee the improvement in both coherence and long-term stability of eNe qubits.
@@ Methods
The resonator, electrode, and on-chip filter were patterned with electron beam lithography followed by reactive-ion etching. See details of the TiN film, resonator, and on-chip filter characterization in the Supplementary I. The chip was mounted on a customized printed circuit board within a vacuum-sealed copper cell, which provides direct current (DC) and microwave (MW) interfaces. On the top lid of the cell, a gas filling line was attached to deposit neon at cryogenic temperature, and a tungsten filament was used as the electron source. The cell is mounted on the mixing chamber (MXC) plate of a dilution refrigerator. A total attenuation of 60 dB was applied on the MW input line with infrared filters (QMC-CRYOIRF-004). The MW output line was equipped with cryogenic isolators (LNF-ISISC4_12A) at MXC plate, followed by a high electron mobility transistor (LNF-LNC4_8C) at 4K plate and room temperature amplification. All DC connections were filtered with thermocoaxes cable, LC filters (Mini-Circuits RLP-30+), and homemade low-pass filters with 150 Hz cutoff. Qubit spectroscopic measurements were conducted with a vector network analyzer (Keysight N5222B) and a signal generator (Anritsu MG3692C). Time-domain pulse measurements were conducted with Quantum Machine OPX+ and Octave. The DC gates were applied with QDevil QDAC-II.
@@
Neon is filled with the following procedure. The fridge is warmed up from its base temperature with a heater mounted on the 4 K plate. At this moment, the helium mixture circulation is turned off and all mixture has been collected, while the pulse tube is still on. The heater power is set to a value such that it creates a temperature gradient from 27 K at the 4 K plate to about 25 K at the MXC plate. At such condition, the neon gas is filled and deposit onto the device chip in its liquid phase. After filling, the heater is turned off and whole fridge is cooled down again to let the liquid neon freeze into solid. During the cool down, we further anneal the neon film at 10 K for 1 hour.
@@
We use the dynamical decoupling technique with Carr-Purcell-Meiboom-Gill (CPMG) pulse sequences to study the spectral distribution of high-frequency noise affecting the eNe qubits [37, 38]. In our sequence, $N$ refocusing $Y_{\pi}$ pulses are applied between two $X_{\pi/2}$ pulses with identical separation $\tau/N$ between two pulses. Under a CPMG sequence, qubits’ coherence decaying follows $P_{e}(N,\tau)=P_{0}+a\cdot\text{exp}\left(-\tau/2T_{1}\right)\text{exp}\left(-\chi_{P}\right)\text{exp}\left(-\chi_{N}(\tau)\right)$, accounting for pure dephasing $-\chi_{N}$, energy relaxation $-\tau/2T_{1}$, and decay during driving pulses $-\chi_{P}$ (ref. [37]). We define $T_{2}^{\text{CPMG}}$ as the time when the qubit’s coherence decays by a factor of $1/e$ due to energy relaxation and pure dephasing. Increasing the number of $Y_{\pi}$ pulses reduces the time the qubit is exposed to noise before each refocusing, thereby extending the coherence time. Under Gaussian noise assumption [37, 38], a CPMG sequence with $N$ refocusing pulses and total delay time of $\tau$ imposes a filter function $g_{N}(\omega,\tau)$ to the qubit noise caused by source $\lambda$ with spectral distribution $S_{\lambda}(\omega)$, which determines the qubit’s dephasing $\chi_{N}(\tau)$:
@@
To investigate the low-frequency dynamics of qubit decoherence near the charge sweet-spot, we repeatedly conducted Ramsey measurements for 128 iterations, with each recording taking 33 seconds [21, 20]. As shown in Fig. Noise-resilient host for electron qubit operation up to 0.4 Kd, we extracted the qubit frequency drift through a fast Fourier transform (FFT) analysis of the measured Ramsey fringes. During the measurement, we also observed more drastic qubit frequency fluctuations on the scale of 10 – 100 MHz, corresponding to 0.14 – 0.44 mV in $\Delta V_{\text{res}}$. Sometimes we even observed the disappear of qubit signatures in the usual gate scanning range. These qubit frequency “jumps” usually occurred less than 0.5 times per hour. We attribute these large fluctuations to charge rearrangements of nearby weakly trapped electrons on neon [52]. For all of presented measurements at various bias points, we ensured that the qubit remained relatively stable and free from such significant fluctuations during data collection.
@@ Acknowledgements.
This work was performed at the Center for Nanoscale Materials, a U.S. Department of Energy Office of Science User Facility, supported by the U.S. DOE, Office of Basic Energy Sciences, under Contract No. DE-AC02-06CH11357. D. J., X. H., and X. L. acknowledge support from Argonne National Laboratory Directed Research and Development (LDRD) program. D. J. and Y. W. acknowledge support from the Air Force Office of Scientific Research (AFOSR) under Award No. FA9550-23-1-0636. D. J. and W. G. acknowledge support from the National Science Foundation (NSF) under Award No. OSI-2426768. D. J., X. Zhou, and Y. H. acknowledge support from the Julian Schwinger Foundation for Physics Research. D. J. acknowledges support from the Department of Energy (DOE) under Award No. DE-SC0025542. B. D. acknowledges support from the NSF under Award No. DMR-1906003. C. S. W. and X. L. acknowledge support from Q-NEXT, one of the US DOE Office of Science National Quantum Information Science Research Centers. W. G. acknowledges support from the Gordon and Betty Moore Foundation through Grant DOI 10.37807/gbmf11567 and the National High Magnetic Field Laboratory at Florida State University, which is supported by the NSF Cooperative Agreement No. DMR-2128556 and the state of Florida. X. Zhang and Y. H. acknowledge support from the Office of Naval Research (ONR) Young Investigator Program (YIP) program under Award No. N00014-23-1-2144. X. H. acknowledges support from France and Chicago Collaborating in the Sciences (FACCTS) program. This work was partially supported by the University of Chicago Materials Research Science and Engineering Center, which is funded by the NSF under award number DMR-2011854. This work made use of the Pritzker Nanofabrication Facility of the Institute for Molecular Engineering at the University of Chicago, which receives support from SHyNE, a node of the NSF National Nanotechnology Coordinated Infrastructure (NSF NNCI-1542205). The authors thank David I. Schuster and Amir Yacoby for helpful discussions.
Author contributions X. L., X. Zhou, and D. J. devised the experiment. X. L. and C. S. W. conducted the experiment. X. L. designed and fabricated the device. X. L., C. S. W., and B. D. analyzed the data. Y. H., X. Zhang and X. H. supported the experimental measurement. Y. W. and W. G. supported the theoretical understanding. D. J. conceived the idea and led the project. X. L., C. S. W., X. Zhou, and D. J. wrote the original manuscript. All authors contributed to the work.

### v2

@@
A separate measurement on a different qubit Q2, presented in Supplementary V, also reveals strikingly thermal behavior in $T_{1}$ and thermal population but with a different $T_{1}$ in the low-temperature limit. It suggests a picture in which individual eNe qubits couple differently to their environment (perhaps through phonons [14, 9]), which sets the low temperature $T_{1}$. This motivates further studies of the microscopic limitations of $T_{1}$. At these qubit frequencies, $T_{1}$ drops to only about half the low-temperature value at 200 mK, facilitating operation at higher temperatures. In contrast, electron qubits in semiconductor materials are susceptible to phonon effects [25, 49], which can significantly degrade $T_{1}$ at elevated temperatures. Given these favorable scalings, we can also anticipate further improvement of the high-temperature performance of eNe qubits by engineering the electron trapping and detection structure to support charge- or spin-qubit operations at higher frequencies [24].
@@
The temperature dependence of the coherence data, however, remains complex. $T_{2}^{*}$ and $T_{2}^{\rm{echo}}$ show that quasi-static and high-frequency noise components behave differently with temperature, as shown in Fig. Noise-resilient solid host for electron qubits above 100 mKb. With increasing temperature, $T_{2}^{*}\propto T^{-0.74}$, indicating the quasi-static noise power scales almost linearly against temperature. Below 50 mK, $T_{2}^{\rm{echo}}$ approaches $2T_{1}$, suggesting that the single echo pulse still effectively mitigates the quasi-static noise that contributes to most of the dephasing. However, starting from 75 mK, $T_{2}^{\rm{echo}}$ scales with temperature as $\propto T^{-1.1}$, and quickly degrades from $2T_{1}$, indicating the rise of non-quasistatic dephasing noises at higher temperatures [23]. Despite that, Q1 maintains $>$ 1 µs $T_{2}^{\rm{echo}}$ at elevated temperatures up to 400 mK with qubit frequency only at similar-to\,$\sim$ 5 GHz, showing eNe’s thermal-resilient coherence performance and potential for high-temperature operation.
@@
Fig. 5. : Temperature-dependent coherence of eNe qubit Q1 at charge sweet-spot. a, $T_{1}$ (blue dots, data) versus mixing chamber (MXC) temperature. The solid curve represents the predicted decay rate $T_{1}(T)=T_{1}(T=0)\cdot\text{tanh}(\hbar\omega_{\rm{q}}/2k_{B}T)$ where we use the measured value at 10 mK for $T_{1}(T=0)$. Extracted electron temperature (gray diamonds, data) versus MXC temperature based on the measured thermal population in Supplementary IV. b, Decoherence time, $T_{2}^{*}$ (orange dots) and $T_{2}^{\rm{echo}}$ (green dots) versus the MXC temperature, with power-law fitting. Gray dots and curves show $2T_{1}$ for comparison. c, Extracted pure dephasing time $T_{\phi}$ as a function of the MXC temperature. Red dashed curve: Parameter-free calculation of the resonator-induced dephasing based on Q1 properties.
@@
We then extract the relation between the pure dephasing time ($T_{\phi}$) and MXC temperature, following the coherence decay formula described in the Methods. As shown in Fig. Noise-resilient solid host for electron qubits above 100 mKc, $T_{\phi}$ begins to decrease significantly when the MXC temperature exceeds 100 mK. In this range, the measured data matches well with a parameter-free model accounting for the effect of resonator thermal photon on the qubit dephasing [24, 50]:
@@
where $\kappa$ is the resonator decay rate, and $\chi$ is the resonator’s dispersive shift (see Supplementary IV). $n_{\text{th}}$ represents the resonator thermal photon population, equal to $1/(e^{hf_{\rm{r}}/k_{\text{B}}T}-1)$. At higher temperatures, qubit coherence is degraded primarily due to thermal effects. Nevertheless, the data’s deviation from the model below 100 mK suggests the presence of other low-temperature dephasing mechanisms. Separate measurements on Q2 (Supplementary V) give similar results.
@@
The inconsistent spectral properties of eNe qubits indicate that the local neon profile is crucial in electron-trapping mechanisms. Variations in the neon surface profile, caused by liquid-phase solidification and substrate roughness, introduce randomness in the energy potential landscape that hosts the electron qubits. Additionally, the relative position of trapped electrons to nearby electrodes determines the frequency lever-arm and the qubit’s sensitivity to voltage fluctuations. To enhance consistency and control over individual qubit properties, refined neon growth methods, and gate-defined electron trapping mechanisms should be developed to mitigate the effects of neon surface randomness.
@@
Finally, the limiting factor for the non-radiative energy relaxation rate of eNe qubits needs further investigation. The non-monotonic variations in $T_{1}$ of qubits with different charge sweet-spot frequencies point to localized non-radiative decay channels. Nearby charges may create a sparse bath of two-level fluctuators weakly coupled to the eNe qubit, whose density determines the transverse noise intensity [53, 51]. Electrical manipulations could be applied to bias or repeal those weakly coupled charges to reduce the relaxation rate [54]. With the development of improved electron loading and trapping methods, we foresee the improvements in both the coherence and long-term stability of eNe qubits.
@@ Methods
The resonator, electrode, and on-chip filter were patterned with electron beam lithography followed by reactive-ion etching. See details of the TiN film, resonator, and on-chip filter characterization in Supplementary I. The experiment setup is similar to the ones in our previous works [6, 7]. The chip was mounted on a customized printed circuit board within a vacuum-sealed copper cell, which provides direct current (DC) and microwave (MW) interfaces. On the top lid of the cell, a gas filling line was attached to deposit neon at cryogenic temperature, and a tungsten filament was used as the electron source. The cell is mounted on the mixing chamber (MXC) plate of a dilution refrigerator. A total attenuation of 60 dB was applied on the cryogenic segment of the MW input line with infrared filters (QMC-CRYOIRF-004). The MW output line was equipped with cryogenic isolators (LNF-ISISC4_12A) at the MXC plate, followed by a high electron mobility transistor (LNF-LNC4_8C) at 4K plate and room temperature amplification. All DC connections were filtered with thermocoaxes cable, LC filters (Mini-Circuits RLP-30+), and homemade low-pass filters with 150 Hz cutoff. Qubit spectroscopic measurements were conducted with a vector network analyzer (Keysight N5222B) and a signal generator (Anritsu MG3692C). Time-domain pulse measurements were conducted with Quantum Machine OPX+ and Octave. The DC gates were applied with QDevil QDAC-II.
@@
Neon is filled with the following procedure. The fridge is warmed up from its base temperature with a heater mounted on the 4 K plate. At this moment, the helium mixture circulation is turned off, and all the mixture has been collected while the pulse tube is still on. The heater power is set to a value such that it creates a temperature gradient from 27 K at the 4 K plate to about 25 K at the MXC plate. Under such conditions, the neon gas is filled and deposited onto the device chip in its liquid phase. After filling, the heater is turned off, and the whole fridge is cooled down again to let the liquid neon freeze into solid. During the cool down, we further anneal the neon film at 10 K for 1 hour.
@@
We use the dynamical decoupling technique with Carr-Purcell-Meiboom-Gill (CPMG) pulse sequences to study the spectral distribution of high-frequency noise affecting the eNe qubits [37, 38]. In the sequence, $N$ refocusing $Y_{\pi}$ pulses are applied between two $X_{\pi/2}$ pulses with identical separation $\tau/N$ between two pulses. Under such sequences, qubits’ coherence decaying follows $P_{e}(N,\tau)=P_{0}+a\cdot\text{exp}\left(-\tau/2T_{1}\right)\text{exp}\left(-\chi_{P}\right)\text{exp}\left(-\chi_{N}(\tau)\right)$, accounting for pure dephasing $-\chi_{N}$, energy relaxation $-\tau/2T_{1}$, and decay during driving pulses $-\chi_{P}$ (ref. [37]). We define $T_{2}^{\text{CPMG}}$ as the time when the qubit’s coherence decays by a factor of $1/e$ due to energy relaxation and pure dephasing. Increasing the number of $Y_{\pi}$ pulses reduces the time the qubit is exposed to noise before each refocusing, thereby extending the coherence time. Under Gaussian noise assumption [37, 38], a CPMG sequence with $N$ refocusing pulses and total delay time of $\tau$ imposes a filter function $g_{N}(\omega,\tau)$ to the qubit noise caused by source $\lambda$ with spectral distribution $S_{\lambda}(\omega)$, which determines the qubit’s dephasing $\chi_{N}(\tau)$:
@@
To investigate the low-frequency dynamics of qubit decoherence near the charge sweet-spot, we repeatedly conducted Ramsey measurements for 128 iterations, with each recording taking 33 seconds [21, 20]. As shown in Fig. Noise-resilient solid host for electron qubits above 100 mKd, we extracted the qubit frequency drift through a fast Fourier transform (FFT) analysis of the measured Ramsey fringes. During the measurement, we also observed more drastic qubit frequency fluctuations on the scale of 10 – 100 MHz, corresponding to 0.14 – 0.44 mV in $\Delta V_{\text{res}}$. Sometimes, we even observed the disappearance of qubit signatures in the usual gate scanning range. These qubit frequency ”jumps” usually occur less than 0.5 times per hour. We attribute these large fluctuations to charge rearrangements of nearby weakly trapped electrons on neon [52]. We ensured that the qubit remained relatively stable and free from such significant fluctuations during data collection for all the presented measurements at various bias points.
Data availability
The data that support the findings of this study are available from the corresponding authors upon request. Source data are provided with this paper.
Code availability
The codes used to perform the experiments and to analyse the data in this work are available from the corresponding authors upon request.
@@ Acknowledgements.
Work performed at the Center for Nanoscale Materials, a U.S. Department of Energy Office of Science User Facility, was supported by the U.S. DOE, Office of Basic Energy Sciences, under Contract No. DE-AC02-06CH11357. D. J., X. H., and X. L. acknowledge support from Argonne National Laboratory Directed Research and Development (LDRD) program. D. J. and Y. W. acknowledge support from the Air Force Office of Scientific Research (AFOSR) under Award No. FA9550-23-1-0636. D. J. and W. G. acknowledge support from the National Science Foundation (NSF) under Award No. OSI-2426768. D. J., X. Zhou, and Y. H. acknowledge support from the Julian Schwinger Foundation for Physics Research. D. J. acknowledges support from the Department of Energy (DOE) under Award No. DE-SC0025542. B. D. acknowledges support from the NSF under Award No. DMR-1906003. C. S. W. and X. L. acknowledge support from Q-NEXT, one of the US DOE Office of Science National Quantum Information Science Research Centers. W. G. acknowledges support from the Gordon and Betty Moore Foundation through Grant DOI 10.37807/gbmf11567 and the National High Magnetic Field Laboratory at Florida State University, which is supported by the NSF Cooperative Agreement No. DMR-2128556 and the state of Florida. X. Zhang and Y. H. acknowledge support from the Office of Naval Research (ONR) Young Investigator Program (YIP) program under Award No. N00014-23-1-2144. X. H. acknowledges support from France and Chicago Collaborating in the Sciences (FACCTS) program. This work was partially supported by the University of Chicago Materials Research Science and Engineering Center, which is funded by the NSF under award number DMR-2011854. This work made use of the Pritzker Nanofabrication Facility of the Institute for Molecular Engineering at the University of Chicago, which receives support from SHyNE, a node of the NSF National Nanotechnology Coordinated Infrastructure (NSF NNCI-1542205). The authors thank David I. Schuster and Amir Yacoby for their helpful discussions.
Author contributions X. L., X. Zhou, and D. J. devised the experiment. X. L. and C. S. W. conducted the experiment. X. L. designed and fabricated the device. X. L., C. S. W., and B. D. analyzed the data. Y. H., X. Zhang ,and X. H. supported the experimental measurement. Y. W. and W. G. supported the theoretical understanding. D. J. conceived the idea and led the project. X. L., C. S. W., X. Zhou, and D. J. wrote the original manuscript. All authors contributed to the work.